1. Write down the following reactions that occur in electrolysis:
a. NaNO3 solution using electrode C
b. Melt CuSO4 with Ni electrode
c. Melted HCl with the Ag electrode

2. Electrolysis for 2 hours with a current of 10 A carried out on plating the ferrous metal with silver. If that solution
used AgNO 3 (A r Ag = 108) then the mass of silver that has deposited is...
1
Expert's answer
2020-12-01T09:18:42-0500

a) NaNO3 + 2H2O = Na+ + NO3- + 2H2 + O2.

b) CuSO4 + Ni + H2O = Cu + NiSO4 + O2.

c) HCL + Ag = H2 + Cl2.

2. m= E x I x t/26.8= 108 x 10 x 2 / 26.8 = 80,5970149.